Calculate the internal energy of a system (in kiloJoules) that loses 8.54 kJ while expanding from 1.0 to 24.42 L against a pressure of 35.08 atm. Answer to one decimal space.
Answer #1

The work performed by the system is calculated:

W = P * ΔV = 35.08 atm * (24.42 - 1) L = 815.26 atm * L

We calculate in kilojoules:

W = 815.26 atm * L * (0.101325 kJ / 1 atm * L) = 82.61 kJ

The internal energy change is calculated:

ΔU = Q - W = 8.54 - 82.61 = - 74.07 kJ


If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.